OpenStudy (anonymous):

HELP MATHHH Write an equation for the line that is parallel to the given line and passes through the given point. Write your equation in slope-intercept form. y = 4x-3;(-3,12)

OpenStudy (anonymous):

Since it's parallel, the slope is the same so it'll be 4 so you plug everything into (y-y1)=m(x-x1) (y-(12))=(4)(x-(-3)) y-12=4x+12 then you just solve for y

OpenStudy (anonymous):

okay thank you so much what about this one Write the equation of a line that is perpendicular to the given line and that passes through the given point. Write your equation in point-slope form. y-10 =3/2(x-4);(4,2)

OpenStudy (anonymous):

Okay since it says perpendicular, the slope must be opposite reciprocal so the slope for the equation is 3/2 the opposite reciprocal of 3/2 is -2/3 then you just plug everything in (y-y1)=m(x-x1) and solve for y

OpenStudy (anonymous):

thanks i plugged the numbers in i have the answer, i really need help on this study guide but im tired so can you help me tomorrow. @m.lupe95

OpenStudy (anonymous):

sure like around what time so I can see if I'm able to be on.

OpenStudy (anonymous):

anytime is find what time would you like?

OpenStudy (anonymous):

I'm free from 3-5:30pm

OpenStudy (anonymous):

okay thats great see you then ... thanksssss ( :

OpenStudy (anonymous):

okay :) no problem!

OpenStudy (anonymous):

@m.lupe95

OpenStudy (nikato):

I can help! Or try. @bettylee

OpenStudy (anonymous):

@nikato pleaseee thanks you

OpenStudy (anonymous):

@nikato please helppp

OpenStudy (anonymous):

what do you need help on?

OpenStudy (anonymous):

What is the slope of a line passing through the pair of points (2, 10) and (8, 3)

OpenStudy (anonymous):

Okay you use the slope formula\[\frac{ y2-y1 }{ x2-x1 }\] plug in the coordinates and what do you get?

OpenStudy (anonymous):

yessss how ?

OpenStudy (anonymous):

you're given (2,10) and (8,3) so you plug them in into the slope formula i posted \[\frac{ (10)-(3) }{(2)-(8) }\]

OpenStudy (anonymous):

so do i work it out ?

OpenStudy (anonymous):

yeah and that's your slope what did you get?

OpenStudy (anonymous):

7/6 ?

OpenStudy (anonymous):

Close.. you forgot the negative from (2)-(8).. you get -6

OpenStudy (anonymous):

so its 7-6 ?

OpenStudy (anonymous):

7/-6 and that's your slope :)

OpenStudy (anonymous):

okay thank you can you help me with a few days ?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

what else do you need help on?

OpenStudy (anonymous):

The table shows the number of miles driven over time. Express the relationship between distance and time in simplified form as a unit rate: ___________ mph. Write a statement that interprets the relationship time- 3 5 7 9 distance-171 285 399 513

OpenStudy (anonymous):

okay can I reply back in a little bit?.. I'm in class and I'm going to try to solve it at home

OpenStudy (anonymous):

okay that's fine ( :

OpenStudy (anonymous):

okay I'm home :) may I ask what topic is this so I review it

OpenStudy (anonymous):

well theres not really a topic more like a study guide my teacher just gave us that's due tomorrow and I only have 6 questions out of 17 so its a lot.

OpenStudy (anonymous):

Ah..okay well what math are you taking?

OpenStudy (anonymous):

im taking algebra 1 right now lol

OpenStudy (anonymous):

okay I'm looking through the notes right now :) .. it's been long since I took Algebra and I forgot how to solve this :)

OpenStudy (anonymous):

its okay I have a couple more problems also.. its really confusing and im passing with a 92

OpenStudy (anonymous):

what other problems do you have? maybe I'll remember some

OpenStudy (anonymous):

12) Write each equation in standard form. (Hint: Standard form has three requirements) a.y=-2/5x+7 B. y=-0.5x+0.9

OpenStudy (anonymous):

well for the standard form you use the formula Ax+By=C

OpenStudy (anonymous):

soooo that means ?

OpenStudy (anonymous):

can you log on tomorrow I have to go

OpenStudy (anonymous):

yeah around what time?

OpenStudy (anonymous):

what time is good for you ?

OpenStudy (anonymous):

I'm free from 1-5

OpenStudy (anonymous):

okay perfect see you then and thanks ( :

OpenStudy (anonymous):

okay then :) and no problem!

OpenStudy (anonymous):

@m.lupe95 im hereeee

OpenStudy (anonymous):

I guess were in different time zones haha..I have class right now so Imma be on in about an hour if that's okay

OpenStudy (anonymous):

okay yes I live in ga but okay ill wait for you ( :

OpenStudy (anonymous):

I'm here :)

OpenStudy (anonymous):

@bettylee are you still on?

OpenStudy (anonymous):

@m.lupe95 yes im here now sorry